Stan invested $17,000, part at 18% and part at 4%.
If the total interest at the end of the year is $1,380, how much did he invest at each rate?

Found 2 solutions by checkley71, scott8148:
Answer by checkley71(8403)

.18X+.04(17,000-X)=1,380
.18X+680-.04X=1,380
.14X=1,380-680
.14X=700
X=700/.14
X=5,000 INVESTED @ 18%
17,000-5,000=12,000 INVESTED @ 4%.
PROOF
.18*5,000 +.04*12,000=1,380
900+480=1,380
1,380=1,380

Answer by scott8148(6628)

let x="part @ 18%", so 17000-x="part @ 4%" ... (18%)(x)+(4%)(17000-x)=1380

.18x+680-.04x=1380 ... .14x=700